A company is considering constructing a plant to manufacture a proposed new product. The land costs
$300,000, the building costs $600,000, the equipment costs $250,000, and $100,000 additional working
capital is required. It is expected that the product will result in sales of $750,000 per year for 10 years, at which time the land can be sold for $400,000, the building for $350,000, and the equipment for
$50,000. All of the working capital would be recovered at the EOY 10. The annual expenses for labor,
materials, and all other items are estimated to total $475,000. If the company requires a MARR of 15%
per year on projects of comparable risk, determine if it should invest in the new product line. Use the
AW method.
1
Expert's answer
2021-10-19T14:45:02-0400


Modified B-C ratio = (PW of annual benefits - PW of annual expenses) / (Land + Building + Equipment + Working capital - PW of salvage value)

"= [(750,000 - 475,000) \times P/A(15%, 10)] / [300,000 + 600,000 + 250,000 + 100,000 - (400,000 + 350,000 + 50,000) x P/F(15%, 10)]\\

= (275,000\times 5.0188) / [1,250,000 - (800,000 \times0.2472)]\\

= 1,380,170 / (1,250,000 - 197,760)\\

= 1,380,170 / 1,052,240\\

= 1.31\\"

Since Modified B-C Ratio is greater than 1, Project is acceptable